OpenStudy (anonymous):

State the vertical asymptote of the rational function. f(x) = quantity x minus six times quantity x plus six divided by quantity x squared minus nine.

OpenStudy (jdoe0001):

\(\bf \cfrac{x-6x+6}{x^2-9} \ \ ?\)

OpenStudy (anonymous):

(x-6)(x+6) --------- x2 - 9

OpenStudy (jdoe0001):

well, keep in mind vertical asymptotes for a rational expression occurs at the zeros of the denominator, so long they don't make the numerator 0 so if we set the denominator to 0 \(\bf x^2-9 =0\\ \text{solve for "x", keeping in mind that}\\ (a-b)(a+b) = (a^2-b^2)\)

OpenStudy (jdoe0001):

what would you get on that for "x"?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

is that right??

OpenStudy (jdoe0001):

\(\bf x^2-9 =0\\ x^2-3^2 \implies (x-3)(x+3)\) so the values for "x" are +3 and -3 at that value for "x", the denominator is 0 now let's check those values against the numerator, replace "x" in the numerator for +3 and then for -3, does it give 0?

OpenStudy (jdoe0001):

so, (3-6)(3+6) = -18 (-3-6)(-3+6) = -18 so zeros of the denominator, do not make the numerator 0, thus those are your vertical asymptotes :) x = 3, x = -3

OpenStudy (anonymous):

omg thats so helpful! thank you so much! <33333

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

would you mind helping me with a few more? this chapter is new and nobody else can help really well

OpenStudy (jdoe0001):

is easier to ask in the channel, that way we can all see it and help and revise each other :)

OpenStudy (anonymous):

the channel? :o im new here idk what you mean :(

OpenStudy (jdoe0001):

ohh, where it says, "ask a question"

OpenStudy (jdoe0001):

if we can see it, we can jump in and revise each other too :)

OpenStudy (anonymous):

ok thanks!!